A reheat cycle with two stages of reheating is executed with steam expanding initially from 20 MPa and . The two reheater pressures are 3.8 MPa and 0.97 MPa, and the steam leaves each reheater at . Condensation occurs at . Compute for  for a 1kg/s of steam.


1
Expert's answer
2021-06-28T08:53:25-0400

"P_1=20MPa ; T_1=540^OC \implies h_1=3365.26kJ/kg"

"P_2=3.8MPa ; T_3=540^OC \implies h_2=2801.35kJ/kg ; h_3=3538.30kJ/kg"

"P_4=P_5=0.97MPa ; T_1=540^OC \implies h_4=2779.70kJ/kg; h_5=3565.48 kJ/kg"

"T_6=60^0C \implies h_6=2608.85 kJ/kg"


Heat added "=\dot{m}((h_3-h_2)+(h_5-h_4))"

Heat added "=1((3538.3-2801.35)+(3565.48-2779.70))=1522.7 kJ\/s"


Work "=\dot{m}((h_1-h_2)+(h_3-h_4)+(h_5-h_6))"

Work ="1((3365.26-2801.35)+(3538.3-2779.70)+(3565.48-2608.85))=2279.1 kJ\/s"


Heat rejected = Work - Heat added

Heat rejected = "2279.1-1522.7=756 kJ\/s"